OpenStudy (anonymous):

I'm not sure how to do this, because every time I try it my teacher says it's wrong... Can someone tell me that i'm doing wrong? Part 1: Write an expression for the following real-world situation. If Mr. Jackson earns $1000 for every used car he sells and $2000 for every new car he sells, how much would he make if he sold "u" used cars and "n" new cars? Part 2: Write a real-world situation of your own and a VARIABLE expression to go with it. what is wrong with this?Timmy's mom took out Timmy's friends, if for each adult costs $10 and for each kid $7, how much would it be? 10a + 7k = 17ak?

OpenStudy (anonymous):

1. 1000u+2000n 2. 10a+7k stays 10a+7k

OpenStudy (anonymous):

10a + 7k = Tommy's mom money Let's evaluate your approach let's say a is 4 and k is 2 so 10(4) + 7(2) = (2)(4)7 40+14 = 56 54 = 56 So, it's false and your equation is wrong.. Did i clarify it?

OpenStudy (anonymous):

false!